OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of -3/4 and passes through the point (4, 5).

OpenStudy (anx):

same as before (y-y1) = m (x-x1) (y- 5) = -3/4(x-4) distrubte the -3/4 (y-5) = (-3/4x +3) add 5 y = -3/4x+8